How much do you win for 2 numbers on the lottery?
You will only win with 2 numbers if they also have the mega ball with them. You will win $10.00.
What is the probability of a 33 percent free throw shooter missing 3 in a row?
You want the probability of miss and miss and miss which is .67 * .67 * .67 = 0.301.
Two identical freight cars roll without friction towards each other on a level track one rolls at 2 m s and the other rolls at 1 m s after the cars collide they couple and roll together with a speed?
a black one
What is the probability of rolling the sum of 2 or the sum of 9?
The answer depends on what sort of dice you roll and how many.
What are some probability words?
Some words that can apply to probability are:
maybe; odds; chance(s); if.
Odds of rolling 6 ones with 6 dice?
Probability of rolling a ' 1 ' with a single cube = 1 / 6
Probability of rolling six 1s with 6 cubes =
(1/6) x (1/6) x (1/6) x (1/6) x (1/6) x (1/6) = 1/46,656 = 0.000021433 (rounded) = 0.0021433%
Suppose that 10 fair coins are tossed Find the numbers of ways of obtaining exactly 4 heads?
It is 10C4 = 10!/(6!*4!) = 10*9*8*7/(4*3*2*1) = 210
50%
How many ways can you arrange the letters in the word Mississippi?
Total arrangements are determined by the equation f(n) = n!, where n is the number of letters in the word, and n! is the factorial function, which is n*(n-1) ... *1. This word has 11! total arrangements.
Distinguishable arrangements are determined by the equation f(n) = n!/(c1!*c2! ... *cn!), where the denominator is the product of the factorials of the count of each unique letter in the word.
There is one "m".
There are four "i"s.
There are four "s"s.
There are two "p"s.
So:
11!/(4!4!2!1!) = 39916800/1152 = 34650 distinguishable arrangements
There are 40 ways.
How many cards in deck of skipbo?
There are 162 cards in SkipBo. 144 cards are numbered 1 through 12 and 18 are SkipBo (wild) cards.
What is the relationship between a sample and population is statistics?
Information obtained from the sample can be extrapolated to the whole population using statistics.
What is the probability of drawing a blue card replacing it and then drawing a blue card?
The answer depends on how many cards there are, what colours, how many of each colour, whether or not the cards are in random order. And since you have not bothered to share any of this crucial information, I cannot provide a more useful answer.
The answer depends on how many cards there are, what colours, how many of each colour, whether or not the cards are in random order. And since you have not bothered to share any of this crucial information, I cannot provide a more useful answer.
The answer depends on how many cards there are, what colours, how many of each colour, whether or not the cards are in random order. And since you have not bothered to share any of this crucial information, I cannot provide a more useful answer.
The answer depends on how many cards there are, what colours, how many of each colour, whether or not the cards are in random order. And since you have not bothered to share any of this crucial information, I cannot provide a more useful answer.
What is the probability of winning the lottery depends on how many tickets you buy?
It is a true statement. If you buy them all, the probability of your winning is 1!
It is a true statement. If you buy them all, the probability of your winning is 1!
It is a true statement. If you buy them all, the probability of your winning is 1!
It is a true statement. If you buy them all, the probability of your winning is 1!
What are 3 probabilities of getting heads when you flip a quarter?
There is the probability of 1/2 if it is a fair coin.
There is the probability of 1 if it is a double-headed coin.
There is the probability of 0 if it is a double-tailed coin.
What does the expression mutually exclusive mean?
Mutually exclusive means if one thing is true the other must be false and vice versa.
If A is true, B is false. If B is true, A is false.
For instance,
The ball was totally red.
The ball was totally blue.
These are mutually exclusive because the ball can only be one or the other.
The ball was red.
The ball was blue.
These are NOT mutually exclusive because the ball could also be red AND blue.
Straight thumbs, hitch hiker's thumb is recessive...so the F1 generation in this example would be heterozygous and have the dominant (straight thumbed) phenotype.